OpenStudy (anonymous):

Choose the Standard Form of the equation y = −1/4 x − 6.

OpenStudy (anonymous):

Standard equation looks like Ax+By=C. So rearrange y = −1/4 x − 6, where x is the first variable, y is the second and it is all equal to a constant.---> (1/4)x+y=-6

OpenStudy (anonymous):

y = −(1/4) x − 6. multiply both sides by 4 give u 4y=-x -24 correct?